(vi) 6x + 3y = 6xy

2x + 4y = 5xy

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Solution: 6x + 3y = 6xy 6/y + 3/x = 6 Let 1/x = m and 1/y = n => 6n +3m = 6 =>3m + 6n-6 = 0…………………….(i) 2x + 4y = 5xy => 2/y + 4/x = 5 => 2n +4m = 5 => 4m+2n-5 = 0……………………..(ii) 3m + 6n – 6 = 0 4m + 2n – 5 = 0 By cross-multiplication method, we get m/(-30 –(-12)) = n/(-24-(-15)) = 1/(6-24) m/-18 = n/-9 = 1/-18 m/-18 = 1/-18 m = 1 n/-9 = 1/-18 n = 1/2 m = 1 and n = 1/2 m = 1/x = 1 and n = 1/y = 1/2 x = 1 and y = 2 Hence, x = 1 and y = 2